# Compatibility of canonical $\ell$-adic local systems on Shimura varieties

### v1

@@ 4. Shimura Varieties
Set $\mathcal{S}=\mathcal{S}_{K_{0},s}$, our integral model over $\mathcal{O}_{E_{K_{0},s}}[1/N]$. Here $N$ is chosen large enough that the conclusion of Corollary 3.4 holds for $N$, and that [3, Theorem 1.3] holds away from this $N$, i.e. for $v\nmid N$, $\mathcal{S}_{K_{0}}\times_{\mathcal{O}_{E}[1/N]}\mathcal{O}_{E_{v}}$ is an integral canonical model. We have the canonical local systems
@@
which we will show in this section form a compatible system. For places $\lambda|\ell$, $\lambda^{\prime}|\ell^{\prime}$ of $\overline{\mathbb{Q}}$, we also denote $\rho_{\ell}$ (resp. $\rho_{\ell^{\prime}}$) by $\rho_{\lambda}$ (resp. $\rho_{\lambda^{\prime}}$) when viewed as valued in $G(\overline{\mathbb{Q}}_{\lambda})$ (resp. $G(\overline{\mathbb{Q}}_{\lambda^{\prime}})$. Let $v$ be a prime of $E_{K_{0},s}$ not dividing $N\ell\ell^{\prime}$. We form $\rho_{\lambda,v}$, the restriction to the $v$-fiber. Since the monodromy groups of the $\rho_{\lambda,v}$ might not be semisimple, we need a small argument to check that Drinfeld’s results ([7]) apply. As explained in Lemma 2.1, for the results of [7, §6] to apply to give a $\lambda^{\prime}$-companion, it suffices to check that the eigenvalues of $r(\rho_{\lambda,v})(\mathrm{Frob}_{x})$ are plain of characteristic $p$ for all $x\in|\mathcal{S}_{\kappa(v)}|$ and representations $r$ of the algebraic monodromy group $G_{\lambda,v}\subset G$ of $\rho_{\lambda,v}$. To check this, it suffices to check that the composite $\pi_{1}(\mathcal{S}_{v})\to G/G^{\mathrm{der}}$ is plain of characteristic $p$, but this follows from the canonical model property on the zero-dimensional Shimura variety associated to $G/G^{\mathrm{der}}$, and in particular is a special case of Lemma 4.1 below. Thus a strong $\lambda^{\prime}$-adic companion $\rho_{\lambda\leadsto\lambda^{\prime},v}$ exists.
Lemma 4.1.
@@
Proof.
This is a standard calculation, but we give some details particularly since we have no reference for the crystalline compatibility. Let $\rho_{s}$ be the specialization of $\rho_{K_{0}}$ along $s$ and likewise $\rho_{\ell,s}$ for its $G(\mathbb{Q}_{\ell})$ projection for all $\ell$. By [16, Lemma 3.2], for any $\alpha\in\mathbb{A}_{E(sK_{0})}^{\times}$ mapping to $\sigma\in\Gamma^{\mathrm{ab}}_{E(sK_{0})}$, we have
$q_{\sigma}a\rho_{s}(\sigma)=r_{x}(\alpha)a,$
i.e., $\rho_{s}(\sigma)=a^{-1}q_{\sigma}^{-1}r_{x}(\alpha)a$, for some $q_{\sigma}\in G(\mathbb{Q})_{+}$. For $\sigma\in\Gamma_{E(sK_{0})_{v}}$ and $\ell\neq p$, by $\ell$-projection we obtain $\rho_{\ell,s}(\sigma)=a_{\ell}^{-1}q_{\sigma}^{-1}a_{\ell}$; since the special point is ¯$\overline{\mathbb{Z}}[1/N]$-rational, for $\sigma\in I_{E(sK_{0})_{v}}$ we get $q_{\sigma}=1$, and for $\sigma=\mathrm{Frob}_{v}$ we find $\rho_{\ell,s}(\mathrm{Frob}_{v})$ is for all $\ell\neq p$ conjugate (in $G(\mathbb{Q}_{\ell})$) to the independent of $\ell$ value $q_{\mathrm{Frob}_{v}}^{-1}$.
Projecting to the $p$-component, for $\sigma\in I_{E(sK_{0})_{v}}$ we thus find $\rho_{p,s}(\sigma)=a_{p}^{-1}r_{x}(\alpha)a_{p}$, where $\alpha\in\mathcal{O}_{v}^{\times}$, and $\rho_{p,s}(\mathrm{Frob}_{v})=a_{p}^{-1}q_{\mathrm{Frob}_{v}}^{-1}r_{x}(\varpi_{v})a_{p}$ for the uniformizer $\varpi_{v}$ mapped to the choice of $\mathrm{Frob}_{v}$. In particular, up to conjugacy the abelian representation $\rho_{p,s}|_{I_{E(sK_{0})_{v}}}$ is given by the algebraic representation $r_{x}$, hence is crystalline. In particular, the functor $\mathrm{Rep}(G_{\mathbb{Q}_{p}})\to\operatorname{F-Isoc}(\kappa(v))$,
$\xi\mapsto D_{\mathrm{cris}}(\xi\circ\rho_{p,s}),$
composed with the canonical (forgetful) fiber functor to $\mathrm{Vect}_{E(sK_{0})_{v}}$, gives by the Tannakian formalism an element $\gamma_{v}\in G(\overline{\mathbb{Q}}_{p})$, well-defined up to inner automorphism (resulting from choosing an isomorphism over some finite extension of $E(sK_{0})_{v}$ between this fiber functor and the canonical fiber functor on $\mathrm{Rep}(G_{\mathbb{Q}_{p}})$), such that for all $\xi\in\mathrm{Rep}(G_{\mathbb{Q}_{p}})$, $\xi(\gamma_{v})$ has the same characteristic polynomial (and even same conjugacy class) as the linearized crystalline Frobenius $\varphi^{[\kappa(v):\mathbb{F}_{p}]}$ acting on $D_{\mathrm{cris}}(\xi\circ\rho_{p,s})$. We check this is the characteristic polynomial of $\xi(\operatorname{inn}(a_{p})(\rho_{p,s}(\mathrm{Frob}_{v}))\cdot r_{x}(\varpi_{v})^{-1})$, i.e. is that of $\xi(q_{\mathrm{Frob}_{v}}^{-1})$.
There is a finite extension $K/\mathbb{Q}_{p}$ such that for any representation $\xi$ of $G_{\mathbb{Q}_{p}}$,
$\xi_{K}\circ\operatorname{inn}(a_{p})(\rho_{p,s}|_{\Gamma_{E(sK_{0})_{v}}})=\bigoplus_{i}\psi_{i}$
for crystalline characters $\psi_{i}\colon\Gamma_{E(sK_{0})_{v}}\to K^{\times}$. Restricting to $I_{E(sK_{0})_{v}}$, this decomposition induces a decomposition $\xi_{K}\circ r_{x}=\bigoplus_{i}\chi_{i}$ for algebraic homomorphisms $\chi_{i}\colon\operatorname{Res}_{E(sK_{0})_{v}/\mathbb{Q}_{p}}\mathbb{G}_{m}\to\operatorname{Res}_{K/\mathbb{Q}_{p}}\mathbb{G}_{m}$ of $\mathbb{Q}_{p}$-tori. A calculation that ultimately rests on Lubin-Tate theory ([5, Proposition B.4], to which strictly speaking we should add the corresponding calculation of crystalline periods for unramified representations) shows that $\varphi^{[\kappa(v):\mathbb{F}_{p}]}$ acts on $D_{\mathrm{cris}}(\xi_{K}\circ\operatorname{inn}(a_{p})(\rho_{p,s}|_{\Gamma_{E(sK_{0})_{v}}}))$ (canonically a direct sum of free rank 1 $E(sK_{0})_{v}\otimes_{\mathbb{Q}_{p}}K$-modules, with $\varphi^{[\kappa(v):\mathbb{Q}_{p}]}$ acting $E(sK_{0})_{v}$-linearly) by
$\bigoplus_{i}\psi_{i}(\operatorname{rec}_{v}(\varpi_{v}))\chi_{i}(\varpi_{v})^{-1}.$
By the calculation of $\rho_{p,s}$, this implies $\varphi^{[\kappa(v):\mathbb{F}_{p}]}$ has the same characteristic polynomial as
$\xi_{K}\left(\operatorname{inn}(a_{p})(\rho_{p,s}(\varpi_{v}))\cdot r_{x}(\varpi_{v})^{-1}\right)=\xi_{K}(q_{\mathrm{Frob}_{v}}^{-1}).$
and thus the same characteristic polynomial as $\xi(q_{\mathrm{Frob}_{v}}^{-1})$.
Since finite-dimensional representations separate semi-simple conjugacy classes, $\rho_{\ell,s}(\mathrm{Frob}_{v})$, $q_{\mathrm{Frob}_{v}^{-1}}$, and $\gamma_{v}$ are equal as elements of $[G{\,\!\sslash\!\,}G](\mathbb{Q})$. ∎
@@
Let $H$ be the algebraic monodromy group of $\rho_{\lambda^{\prime},v}$. Our setup implies that $H$ lies between $G^{\mathrm{der}}$ and $G$ (but may be neither semisimple nor connected). From the known adjoint case ([16, Theorem 3.10]) and the abelian compatibility (Lemma 4.1), after suitable conjugation we may assume $\rho_{\lambda\leadsto\lambda^{\prime},v}=\rho_{\lambda^{\prime},v}\cdot\chi$ for some character (depending on $\lambda$, $\lambda^{\prime}$, and $v$) $\chi\colon\pi_{1}(\mathcal{S}_{\kappa(v)})\to Z_{G^{\mathrm{der}}}(\overline{\mathbb{Q}}_{\lambda^{\prime}})$. We know from Theorem 4.2 that $\rho_{\lambda\leadsto\lambda^{\prime},v}(\mathrm{Frob}_{x})$ and $\rho_{\lambda^{\prime},v}(\mathrm{Frob}_{x})$ are conjugate for closed points $x$ in the Dirichlet density 1 (because Zariski open and dense: see Lemma 4.5) $\mu$-ordinary locus of $\mathcal{S}_{\kappa(v)}$. In particular, for any irreducible representation $R\colon G\to\mathrm{GL}_{N}$, by the Čebotarev density theorem (and Brauer-Nesbitt theorem), there is $g\in\mathrm{GL}_{N}(\overline{\mathbb{Q}}_{\lambda^{\prime}})$ such that
@@
In particular, $\mathrm{tr}(R(\rho_{\lambda^{\prime},v}))=\mathrm{tr}(R(\rho_{\lambda^{\prime},v}))\cdot R(\chi)$, where here we view $R(\chi)$ as a scalar (by Schur’s lemma) in ¯$\overline{\mathbb{Q}}_{\lambda^{\prime}}^{\times}$. Now, either $R(\chi)(\mathrm{Frob}_{x})=1$ for a density 1 set of closed points $x$, in which case $R(\chi)=1$, or there is a positive density set of $x$ such that $\mathrm{tr}(R(\rho_{\lambda^{\prime},v})(\mathrm{Frob}_{x}))$ equals zero. In the latter case, by [20, Theorem 3], there is a connected component $\Phi$ of the algebraic monodromy group $H_{R}$ of $R(\rho_{\lambda^{\prime},v})$ on which the trace vanishes identically. [^10] Compare [18, Proposition 3.4.9]. Strictly speaking, Rajan proves his Theorem for representations of $\pi_{1}(\mathcal{O}_{F}[1/N])$, $F$ a number field, but the same arguments will apply for integral finite type $\mathbb{Z}$-schemes, where we still have the Čebotarev density theorem. For some $z\in Z_{G}$, $\Phi=R(zH^{0})$ ($H$ surjects onto $H_{R}$, $H^{0}$ surjects onto $H_{R}^{0}$, and every element of $H/H^{0}$ is represented by an element of $Z_{G}$ since $H\supset G^{\mathrm{der}}$), and since $R(z)$ is a scalar, $\mathrm{tr}(\Phi)=0$ forces $\mathrm{tr}(R(H^{0}))=0$; this is clearly impossible, since $1\in H^{0}$. We conclude that $R(\chi)=1$. Knowing this for all irreducible representations $R$ of $G$, we conclude that $\chi=1$.
@@
We next address the crystalline compatibility. The proof uses the same principle as the proof of Theorem 4.4, except we need the existence of $p$$-\ell$ companions in our setting, as explained in [9]. By [19, Theorem 7.1], for any place $v|p$ of $E_{K_{0},s}$ not dividing $N$ the $p$-adic local system $\rho_{p,v}\colon\pi_{1}(\mathcal{S}_{(E_{K_{0},s})_{v}})\to G(\mathbb{Q}_{p})$ is crystalline. Our joint work [9] explains how to extract from this an ($G$-) overconvergent $F$-isocrystal, packaged as a $G(\overline{\mathbb{Q}}_{p})$-conjugacy class of representations
@@
compatible (via $D_{\mathrm{cris}}$) with $\rho_{p,v}$. For notational consistency, we will also for a place $\lambda$ of $\overline{\mathbb{Q}}$ above $p$ write these representations as $\rho_{\lambda,v}^{\operatorname{F-Isoc}^{\dagger}}$ and $\rho_{\lambda,v}$. We further know ([9, Proposition 4.3.1]) that the image of $\rho_{\lambda,v}^{\operatorname{F-Isoc}^{\dagger}}$ in $G^{\mathrm{ad}}$ is Zariski-dense, and from Lemma 4.1 we see that for any closed point $x\in\mathcal{S}_{v}$ that lifts to a special point, the linearized crystalline Frobenius of $\rho_{\lambda,v}^{\operatorname{F-Isoc}^{\dagger}}|_{\kappa(x)}$ agrees in $[G{\,\!\sslash\!\,}G](\mathbb{Q})$ with $\rho_{\lambda^{\prime},v}(\mathrm{Frob}_{x})$ ($\lambda^{\prime}$ not above $p$). [^11] Here and in what follows, for $\lambda^{\prime}|p$ there is also a $\lambda^{\prime}$-companion, but since $\rho_{p,v}$ has $\mathbb{Q}_{p}$-coefficients, the “$p$-to-$p$” companion construction is just the identity. This compatibility on special points also implies the compatibility of the abelianizations of $\rho_{\lambda,v}^{\operatorname{F-Isoc}^{\dagger}}$ and $\rho_{\lambda^{\prime},v}$. Combining work of Drinfeld ([7]), Abe ([1]), Abe-Esnault ([2]), and Kedlaya ([10]), $\rho_{\lambda,v}^{\operatorname{F-Isoc}^{\dagger}}$ admits a $\lambda^{\prime}$-companion $\rho_{\lambda\rightsquigarrow\lambda^{\prime},v}$ as explained in [9, §4] whose algebraic monodromy group also contains $G^{\mathrm{der}}$.
@@
We conclude with a refinement of Corollary 4.6 that supplies one of the desiderata in the production of Kottwitz-triples (see the introduction). Let $v|p$ be a place not above $N$ of $E^{\prime}$, and let $x\in\mathcal{S}_{v}$ be a closed point. By construction, for all $\lambda^{\prime}$ not above $p$, $\rho_{\lambda^{\prime}}(\mathrm{Frob}_{x})$ lies in $G(\mathbb{Q}_{\ell^{\prime}})$, is semisimple by [3, Theorem 7.1], and we have seen this is an independent of $\lambda^{\prime}$ conjugacy class in $[G{\,\!\sslash\!\,}G](\mathbb{Q})$. For $\lambda|p$, let us write $\rho_{\lambda,v}^{\operatorname{F-Isoc}^{\dagger}}(\mathrm{Frob}_{x})\in G(\overline{\mathbb{Q}}_{p})$ for the linearized crystalline Frobenius of $\rho_{\lambda,v}^{\operatorname{F-Isoc}^{\dagger}}|_{\kappa(x)}$. Corollary 4.6 tells us the semisimple part $\gamma_{x}$ of $\rho_{\lambda,v}^{\operatorname{F-Isoc}^{\dagger}}(\mathrm{Frob}_{x})\in G(\overline{\mathbb{Q}}_{p})$ is $G(\overline{\mathbb{Q}}_{p})$-conjugate to an element of $G(\overline{\mathbb{Q}})$ whose conjugacy class is $\mathbb{Q}$-rational (induced by $\rho_{\lambda^{\prime}}(\mathrm{Frob}_{x})$). Note that we only at present know that $\rho^{\operatorname{F-Isoc}^{\dagger}}_{\lambda,v}(\mathrm{Frob}_{x})$ is semisimple for $x$ in the $\mu$-ordinary locus, since $x$ then lifts to a special point in characteristic zero.
@@
Recall that $K_{p}=G_{\mathbb{Z}_{p}}(\mathbb{Z}_{p})$ is a hyperspecial level structure, where we write $G_{\mathbb{Z}_{p}}$ for an extension of $G$ to a reductive model over $\mathbb{Z}_{p}$. Choose a faithful representation (over $\mathbb{Z}_{p}$) $L$ of $G_{\mathbb{Z}_{p}}$ and set $V=L\otimes_{\mathbb{Z}_{p}}\mathbb{Q}_{p}$. Any choice of lift $\tilde{x}\in\mathcal{S}(W(\kappa(x)))$ induces on $L$ the structure of a crystalline $\pi_{1}(K_{x})$-representation, with underlying $F$-isocrystal independent of the choice of lift $\tilde{x}$. By [12, Proposition 1.3.2], there is a finite set of tensors $(s_{\alpha})\subset L^{\otimes}$ (the direct sum of all $\mathbb{Z}_{p}$-modules formed from $L$ using duals, tensor products, and symmetric and exterior powers) such that $G_{\mathbb{Z}_{p}}$ is the subgroup of $\mathrm{GL}(L)$ fixing (pointwise) all $s_{\alpha}$. Let $\mathfrak{M}$ be Kisin’s fully faithful tensor functor ([11], [12, §1.2])
@@
The tensors $s_{\alpha}\in L^{\otimes}$ transport via $\mathfrak{M}$ to tensors $(\tilde{s}_{\alpha})\subset\mathfrak{M}(L)^{\otimes}$; Galois-invariance of the $s_{\alpha}$ yields $\varphi$-invariance of the $\tilde{s}_{\alpha}$, and then [12, Corollary 1.3.5] shows that the $(\tilde{s}_{\alpha})$ define a reductive group $G_{\mathfrak{M}}\subset\mathrm{GL}(\mathfrak{M}(L))$ isomorphic to $G_{\mathbb{Z}_{p}}\times_{\mathbb{Z}_{p}}\mathfrak{S}$. More precisely, the isomorphism arises from trivializing the $G_{\mathbb{Z}_{p}}\times_{\mathbb{Z}_{p}}\mathfrak{S}$-torsor $\mathcal{P}:=\underline{\mathrm{Isom}}_{\mathfrak{S}}(L\otimes_{\mathbb{Z}_{p}}\mathfrak{S},\mathfrak{M}(L))$ of linear isomorphisms mapping $s_{\alpha}$ to $\tilde{s}_{\alpha}$. Specializing along $u\mapsto 0$, we obtain tensors $\tilde{s}_{\alpha,0}\in(\mathfrak{M}(L)/u\mathfrak{M}(L))^{\otimes}$ and a corresponding trivialization.
@@
that carries $s_{\alpha}$ to $s_{\alpha}$, since the tensors $\tilde{s}_{\alpha}$ and thus $\tilde{s}_{\alpha,0}$ are $\varphi$-invariant. The composite $\Phi$ above can therefore be uniquely expressed as $b\sigma$ for some $b\in G(K_{x})$. Changing the choice of trivialization of $\mathcal{P}\pmod{u}$ amounts to replacing $\gamma$ by $\gamma s$, $s\in G(W(\kappa(x)))$, which changes $\Phi$ to $s^{-1}\Phi s$ and changes $b$ to $s^{-1}b\sigma(s)$. Finally, $\Phi^{d_{x}}=b\sigma(b)\cdots\sigma^{d_{x}-1}(b)$ is an element of $G(K_{x})$ in the same $G(\overline{\mathbb{Q}}_{p})$-conjugacy class as $\rho^{\operatorname{F-Isoc}^{\dagger}}_{\lambda,v}(\mathrm{Frob}_{x})$, since $\mathfrak{M}(L)/u\mathfrak{M}(L)[1/p]$ is canonically $D_{\mathrm{cris}}(V)$ as $\varphi$-isocrystals ([12, Theorem 1.2.1]). ∎

### v2

@@ 4. Shimura Varieties
Set $\mathcal{S}=\mathcal{S}_{K_{0},s}$, our integral model over $\mathcal{O}_{E_{K_{0},s}}[1/N]$. Here $N$ is chosen large enough that the conclusion of Corollary 3.4 holds for $N$, and that [4, Theorem 1.3] holds away from this $N$, i.e. for $v\nmid N$, $\mathcal{S}_{K_{0}}\times_{\mathcal{O}_{E}[1/N]}\mathcal{O}_{E_{v}}$ is an integral canonical model. We have the canonical local systems
@@
which we will show in this section form a compatible system. For places $\lambda|\ell$, $\lambda^{\prime}|\ell^{\prime}$ of $\overline{\mathbb{Q}}$, we also denote $\rho_{\ell}$ (resp. $\rho_{\ell^{\prime}}$) by $\rho_{\lambda}$ (resp. $\rho_{\lambda^{\prime}}$) when viewed as valued in $G(\overline{\mathbb{Q}}_{\lambda})$ (resp. $G(\overline{\mathbb{Q}}_{\lambda^{\prime}})$. Let $v$ be a prime of $E_{K_{0},s}$ not dividing $N\ell\ell^{\prime}$. We form $\rho_{\lambda,v}$, the restriction to the $v$-fiber. Since the monodromy groups of the $\rho_{\lambda,v}$ might not be semisimple, we need a small argument to check that Drinfeld’s results ([8]) apply. As explained in Corollary 2.5, for the results of [8, §6] to apply to give a $\lambda^{\prime}$-companion, it suffices to check that the eigenvalues of $r(\rho_{\lambda,v})(\mathrm{Frob}_{x})$ are plain of characteristic $p$ for all $x\in|\mathcal{S}_{\kappa(v)}|$ and representations $r$ of the algebraic monodromy group $G_{\lambda,v}\subset G$ of $\rho_{\lambda,v}$. To check this, it suffices to check that the composite $\pi_{1}(\mathcal{S}_{v})\to G/G^{\mathrm{der}}$ is plain of characteristic $p$ (by [5, Proposition 4.3] or Lemma 2.2), but this follows from the canonical model property on the zero-dimensional Shimura variety associated to $G/G^{\mathrm{der}}$, and in particular is a special case of Lemma 4.1 below. Thus a strong $\lambda^{\prime}$-adic companion $\rho_{\lambda\leadsto\lambda^{\prime},v}$ exists.
We recall the following $-\ell$ calculation for special points:
Lemma 4.1 (Lemma 5.5 of [10]).
@@
Let $H$ be the algebraic monodromy group of $\rho_{\lambda^{\prime},v}$. Our setup implies that $H$ lies between $G^{\mathrm{der}}$ and $G$ (but may be neither semisimple nor connected). From the known adjoint case ([17, Theorem 3.10]) and the abelian compatibility (Lemma 4.1), after suitable conjugation we may assume $\rho_{\lambda\leadsto\lambda^{\prime},v}=\rho_{\lambda^{\prime},v}\cdot\chi$ for some character (depending on $\lambda$, $\lambda^{\prime}$, and $v$) $\chi\colon\pi_{1}(\mathcal{S}_{\kappa(v)})\to Z_{G^{\mathrm{der}}}(\overline{\mathbb{Q}}_{\lambda^{\prime}})$. We know from Theorem 4.2 that $\rho_{\lambda\leadsto\lambda^{\prime},v}(\mathrm{Frob}_{x})$ and $\rho_{\lambda^{\prime},v}(\mathrm{Frob}_{x})$ are conjugate for closed points $x$ in the Dirichlet density 1 (because Zariski open and dense: see Lemma 4.5) $\mu$-ordinary locus of $\mathcal{S}_{\kappa(v)}$. In particular, for any irreducible representation $R\colon G\to\mathrm{GL}_{N}$, by the Čebotarev density theorem (and Brauer-Nesbitt theorem), there is $g\in\mathrm{GL}_{N}(\overline{\mathbb{Q}}_{\lambda^{\prime}})$ such that
@@
In particular, $\mathrm{tr}(R(\rho_{\lambda^{\prime},v}))=\mathrm{tr}(R(\rho_{\lambda^{\prime},v}))\cdot R(\chi)$, where here we view $R(\chi)$ as a scalar (by Schur’s lemma) in ¯$\overline{\mathbb{Q}}_{\lambda^{\prime}}^{\times}$. Now, either $R(\chi)(\mathrm{Frob}_{x})=1$ for a density 1 set of closed points $x$, in which case $R(\chi)=1$, or there is a positive density set of $x$ such that $\mathrm{tr}(R(\rho_{\lambda^{\prime},v})(\mathrm{Frob}_{x}))$ equals zero. In the latter case, by [21, Theorem 3], there is a connected component $\Phi$ of the algebraic monodromy group $H_{R}$ of $R(\rho_{\lambda^{\prime},v})$ on which the trace vanishes identically. [^8] Compare [19, Proposition 3.4.9]. Strictly speaking, Rajan proves his Theorem for representations of $\pi_{1}(\mathcal{O}_{F}[1/N])$, $F$ a number field, but the same arguments will apply for integral finite type $\mathbb{Z}$-schemes, where we still have the Čebotarev density theorem. For some $z\in Z_{G}$, $\Phi=R(zH^{0})$ ($H$ surjects onto $H_{R}$, $H^{0}$ surjects onto $H_{R}^{0}$, and every element of $H/H^{0}$ is represented by an element of $Z_{G}$ since $H\supset G^{\mathrm{der}}$), and since $R(z)$ is a scalar, $\mathrm{tr}(\Phi)=0$ forces $\mathrm{tr}(R(H^{0}))=0$; this is clearly impossible, since $1\in H^{0}$. We conclude that $R(\chi)=1$. Knowing this for all irreducible representations $R$ of $G$, we conclude that $\chi=1$.
@@
We next address the crystalline compatibility. The proof uses the same principle as the proof of Theorem 4.4, except we need the existence of $p$$-\ell$ companions in our setting, as explained in [10] in the adjoint case and in Corollary 2.5 (or Corollary 2.4) in general. By [20, Theorem 7.1], for any place $v|p$ of $E_{K_{0},s}$ not dividing $N$ the $p$-adic local system $\rho_{p,v}\colon\pi_{1}(\mathcal{S}_{(E_{K_{0},s})_{v}})\to G(\mathbb{Q}_{p})$ is crystalline. Our joint work [10] explains how to extract from this an ($G$-) overconvergent $F$-isocrystal, packaged as a $G(\overline{\mathbb{Q}}_{p})$-conjugacy class of representations
@@
compatible (via $D_{\mathrm{cris}}$) with $\rho_{p,v}$. For notational consistency, we will also for a place $\lambda$ of $\overline{\mathbb{Q}}$ above $p$ write these representations as $\rho_{\lambda,v}^{\operatorname{F-Isoc}^{\dagger}}$ and $\rho_{\lambda,v}$. We further know ([10, Proposition 4.3.1]) that the image of $\rho_{\lambda,v}^{\operatorname{F-Isoc}^{\dagger}}$ in $G^{\mathrm{ad}}$ is Zariski-dense, and from Lemma 4.1 we see that for any closed point $x\in\mathcal{S}_{v}$ that lifts to a special point, the linearized crystalline Frobenius of $\rho_{\lambda,v}^{\operatorname{F-Isoc}^{\dagger}}|_{\kappa(x)}$ agrees in $[G{\,\!\sslash\!\,}G](\mathbb{Q})$ with $\rho_{\lambda^{\prime},v}(\mathrm{Frob}_{x})$ ($\lambda^{\prime}$ not above $p$). [^9] Here and in what follows, for $\lambda^{\prime}|p$ there is also a $\lambda^{\prime}$-companion, but since $\rho_{p,v}$ has $\mathbb{Q}_{p}$-coefficients, the “$p$-to-$p$” companion construction is just the identity. This compatibility on special points also implies the compatibility of the abelianizations of $\rho_{\lambda,v}^{\operatorname{F-Isoc}^{\dagger}}$ and $\rho_{\lambda^{\prime},v}$. Combining work of Drinfeld ([8]), Abe ([1]), Abe-Esnault ([3]), and Kedlaya ([11]), we have seen in Corollary 2.5 that $\rho_{\lambda,v}^{\operatorname{F-Isoc}^{\dagger}}$ admits a $\lambda^{\prime}$-companion $\rho_{\lambda\rightsquigarrow\lambda^{\prime},v}$ whose algebraic monodromy group also contains $G^{\mathrm{der}}$.
@@
We conclude with a refinement of Corollary 4.6 that supplies one of the desiderata in the production of Kottwitz-triples (see the introduction). Let $v|p$ be a place not above $N$ of $E^{\prime}$, and let $x\in\mathcal{S}_{v}$ be a closed point. By construction, for all $\lambda^{\prime}$ not above $p$, $\rho_{\lambda^{\prime}}(\mathrm{Frob}_{x})$ lies in $G(\mathbb{Q}_{\ell^{\prime}})$, is semisimple by [4, Theorem 7.1], and we have seen this is an independent of $\lambda^{\prime}$ conjugacy class in $[G{\,\!\sslash\!\,}G](\mathbb{Q})$. For $\lambda|p$, let us write $\rho_{\lambda,v}^{\operatorname{F-Isoc}^{\dagger}}(\mathrm{Frob}_{x})\in G(\overline{\mathbb{Q}}_{p})$ for the linearized crystalline Frobenius of $\rho_{\lambda,v}^{\operatorname{F-Isoc}^{\dagger}}|_{\kappa(x)}$. Corollary 4.6 tells us the semisimple part $\gamma_{x}$ of $\rho_{\lambda,v}^{\operatorname{F-Isoc}^{\dagger}}(\mathrm{Frob}_{x})\in G(\overline{\mathbb{Q}}_{p})$ is $G(\overline{\mathbb{Q}}_{p})$-conjugate to an element of $G(\overline{\mathbb{Q}})$ whose conjugacy class is $\mathbb{Q}$-rational (induced by $\rho_{\lambda^{\prime}}(\mathrm{Frob}_{x})$). Note that we only at present know that $\rho^{\operatorname{F-Isoc}^{\dagger}}_{\lambda,v}(\mathrm{Frob}_{x})$ is semisimple for $x$ in the $\mu$-ordinary locus, since $x$ then lifts to a special point in characteristic zero.
@@
Recall that $K_{p}=G_{\mathbb{Z}_{p}}(\mathbb{Z}_{p})$ is a hyperspecial level structure, where we write $G_{\mathbb{Z}_{p}}$ for an extension of $G$ to a reductive model over $\mathbb{Z}_{p}$. Choose a faithful representation (over $\mathbb{Z}_{p}$) $L$ of $G_{\mathbb{Z}_{p}}$ and set $V=L\otimes_{\mathbb{Z}_{p}}\mathbb{Q}_{p}$. Any choice of lift $\tilde{x}\in\mathcal{S}(W(\kappa(x)))$ induces on $L$ the structure of a crystalline $\pi_{1}(K_{x})$-representation, with underlying $F$-isocrystal independent of the choice of lift $\tilde{x}$. By [13, Proposition 1.3.2], there is a finite set of tensors $(s_{\alpha})\subset L^{\otimes}$ (the direct sum of all $\mathbb{Z}_{p}$-modules formed from $L$ using duals, tensor products, and symmetric and exterior powers) such that $G_{\mathbb{Z}_{p}}$ is the subgroup of $\mathrm{GL}(L)$ fixing (pointwise) all $s_{\alpha}$. Let $\mathfrak{M}$ be Kisin’s fully faithful tensor functor ([12], [13, §1.2])
@@
The tensors $s_{\alpha}\in L^{\otimes}$ transport via $\mathfrak{M}$ to tensors $(\tilde{s}_{\alpha})\subset\mathfrak{M}(L)^{\otimes}$; Galois-invariance of the $s_{\alpha}$ yields $\varphi$-invariance of the $\tilde{s}_{\alpha}$, and then [13, Corollary 1.3.5] shows that the $(\tilde{s}_{\alpha})$ define a reductive group $G_{\mathfrak{M}}\subset\mathrm{GL}(\mathfrak{M}(L))$ isomorphic to $G_{\mathbb{Z}_{p}}\times_{\mathbb{Z}_{p}}\mathfrak{S}$. More precisely, the isomorphism arises from trivializing the $G_{\mathbb{Z}_{p}}\times_{\mathbb{Z}_{p}}\mathfrak{S}$-torsor $\mathcal{P}:=\underline{\mathrm{Isom}}_{\mathfrak{S}}(L\otimes_{\mathbb{Z}_{p}}\mathfrak{S},\mathfrak{M}(L))$ of linear isomorphisms mapping $s_{\alpha}$ to $\tilde{s}_{\alpha}$. Specializing along $u\mapsto 0$, we obtain tensors $\tilde{s}_{\alpha,0}\in(\mathfrak{M}(L)/u\mathfrak{M}(L))^{\otimes}$ and a corresponding trivialization.
@@
that carries $s_{\alpha}$ to $s_{\alpha}$, since the tensors $\tilde{s}_{\alpha}$ and thus $\tilde{s}_{\alpha,0}$ are $\varphi$-invariant. The composite $\Phi$ above can therefore be uniquely expressed as $b\sigma$ for some $b\in G(K_{x})$. Changing the choice of trivialization of $\mathcal{P}\pmod{u}$ amounts to replacing $\gamma$ by $\gamma s$, $s\in G(W(\kappa(x)))$, which changes $\Phi$ to $s^{-1}\Phi s$ and changes $b$ to $s^{-1}b\sigma(s)$. Finally, $\Phi^{d_{x}}=b\sigma(b)\cdots\sigma^{d_{x}-1}(b)$ is an element of $G(K_{x})$ in the same $G(\overline{\mathbb{Q}}_{p})$-conjugacy class as $\rho^{\operatorname{F-Isoc}^{\dagger}}_{\lambda,v}(\mathrm{Frob}_{x})$, since $\mathfrak{M}(L)/u\mathfrak{M}(L)[1/p]$ is canonically $D_{\mathrm{cris}}(V)$ as $\varphi$-isocrystals ([13, Theorem 1.2.1]). ∎